DIRECT-POTENTIAL-FIT DETERMINATION OF AN ACCURATE ANALYTICAL POTENTIAL FOR THE B^{1}Π_{u} “BARRIER” STATE OF Li_{2}

\underline{YIYE HUANG} and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

The potential energy curve for the B^{1}Π_{u} state of Li_{2} is known to have an unusual shape, in that its long-range interaction is dominated by a repulsive term which gives this potential a rotationless barrier which protrudes above its energy asymptote. We have used program DSPotFit^{a} in a direct fit of available spectroscopic data for this state, including the observed Λ– doubling splittings and tunneling predissociation line widths, to determine an analytic potential energy function plus Born-Oppenheimer breakdown and Λ–doubling radial correction functions. Differences between this approach and a recent direct-potential-fit analysis^{b} using a numerical model potential function will be critically examined.

a R.J. Le Roy, J.Y. Seto and Y. Huang, University of Waterloo Chemical Physics Research Report CP-651 (2001); also available through http://leroy.uwaterloo.ca.

b N. Bouloufa, P. Cacciani, R. Vetter, A. Yiannopoulou, F. Martin and A.J. Ross, J. Chem. Phys. (2001, in press).